P9-12 (similar to) E Question Help The effect of tax rate on WACC K Bell Jewelers wishes to explore the efflect on its cost of capital of the rate at which the company pays taxes. The firm wishes to maintain a capital financing is 12%, and the before tax cost of debt finacing is 6% Calculate the weighted average cost of capital MACC) g en a tax rate of 25% 40 % debt 15% preferred stock, and 45% common stock The cost of financing with retar ed e earnings 18%, the cost of preferred stock The firms WACC is[% (Round to two decrmal places) All parts Clear All arch
Answer #1

WACC = were + wd(1-t)rd + wprp

where, we, wd, wp are the weights of equity, debt and preferred stock

re, rd, rp are the cost of financing of equity debt and preferred stock

t is the tax rate

Given, we = 0.45

wd = 0.40

wp = 0.15

re = 0.18

rd = 0.06

rp = 0.12

t = 0.15

Hence, WACC = 0.45*0.18 + 0.40*(1-0.25)*0.06 + 0.15*0.12 = 0.117

Hence, WACC = 11.70%